OpenStudy (fanduekisses):

How do I find the point slope when I only have a slope and a y-intercept?

OpenStudy (fanduekisses):

for ex: slope= -5, y-intercept=-8

OpenStudy (anonymous):

if you know the slope and the y intercept; you can plug in values for x to get the value for y

OpenStudy (anonymous):

y = -5x - 8

OpenStudy (anonymous):

that's your slope intercept form; do you know the formula for point slope?

OpenStudy (fanduekisses):

omg thanks, that was easy lol, >.<

OpenStudy (anonymous):

You're welcome

OpenStudy (fanduekisses):

Sometimes I get confused :p

OpenStudy (fanduekisses):

oh no wait

OpenStudy (anonymous):

Just to be clear, the equation I gave you isn't the final answer.

OpenStudy (fanduekisses):

I am looking for the point slope

OpenStudy (fanduekisses):

Yeah I just realized lol

OpenStudy (fanduekisses):

sorry

OpenStudy (fanduekisses):

I need to find the x intercept

OpenStudy (fanduekisses):

right?

OpenStudy (anonymous):

to find the x int you set y = 0

OpenStudy (fanduekisses):

ok

OpenStudy (anonymous):

But really you just need to rearrange to slope intercept form I gave you.

OpenStudy (anonymous):

rearrange it to \[(y-y _{1}) = m(x-x_{1})\]

OpenStudy (fanduekisses):

yes,but what do i put for x?

OpenStudy (fanduekisses):

(y-(-8)=-5(x-?)

OpenStudy (anonymous):

Try substitution values in the for x in the slope intercept form; then plug those values into the point slope form. m or in this case, -5, does not change

OpenStudy (fanduekisses):

I'm confused.